Recent Log Entries

Visual inspection of 8E02 region

Log Number: 4008252Logbook(s): Body: 

We went down under control access this morning to do a quick inspection of the beamline components around 8E01 and 8E02. A majority of the magnets in the area were last measured in 2014. Using a precision level we checked both quad stacks and saw no real indication of any rotation issues. We did find (4) dipoles rolled clockwise looking downstream about the beam axis. I estimate that this rotation is about 0.05 degrees for all magnets. The dipoles in question are MBY6E03, MBY8E03, MYA8T01 and MYA6T01A. The YA magnets have vacuum chambers with two beam pipes.

The following cavity histories have been updated: 1L11-5

Log Number: 4008247Logbook(s): Tags: Body: 

The following cavity histories have been updated:
Username: tdavis
Date/Time: 06-28-22_10:34:58
RF Status: Good
Tuner Status: Good
Comment: Replaced shorted klystron (SN: Varian #91R3) with good tested klystron (L3 #100165). Recovered the cavity into RF.
Special Instructions:

The OPS DRVHs for the following cavities have been updated: 1L15-3

Log Number: 4008238Logbook(s): Tags: Body: 

The OPS DRVHs for the following cavities have been updated:
Username: michaelm
Date/Time: 06-28-22_10:15:39
RF Status: Good
Tuner Status: Good
Comment: DRVH=13.100 Recovered. Planetary drive tuner; leave in turbo. TDOFF -28 degrees; don't re-zero.
Previous OPS DRVH: 10
Special Instructions:

Box Supplies Turned on and Cycled Once

Log Number: 4008235Logbook(s): Tags: Body: 

The following box supplies were TURNED ON AND CYCLED ONCE by version 2-7 of the Box Supply Cycling Script:
ARC1A Ramp Rate = 8.025 A/s
ARC2A Ramp Rate = 6.987 A/s
ARC3A Ramp Rate = 9.786 A/s
ARC4A Ramp Rate = 11.203 A/s
ARC5A Ramp Rate = 10.304 A/s
ARC6A Ramp Rate = 9.105 A/s
ARC7A Ramp Rate = 10.351 A/s
ARC8A Ramp Rate = 11.815 A/s
ARC9A Ramp Rate = 9.092 A/s
ARCAA Ramp Rate = 12.076 A/s
RSEP8A Ramp Rate = 10.630 A/s
RSEP9A Ramp Rate = 9.380 A/s
XSEP2T Ramp Rate = 5.753 A/s
XSEP4T Ramp Rate = 5.314 A/s
XSEP6T Ramp Rate = 4.989 A/s
XSEP8T Ramp Rate = 17.787 A/s

ioctsbhrp restored

Log Number: 4008231Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2022-06-28 09:34:58 The output from the myRestore execution was as follows: ioctsbhrp pcharp@2022-06-28 09:34 12(12) restored systemEnum@2022-06-28 09:34 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i ioctsbhrp (success) Procedure return status: Success

RF Module Downloads

Log Number: 4008227Logbook(s): Body: COMMENTS: RFCM replaced for PMES unstable.

The following module files were successfully downloaded
nl031e nl031d

6/27 X ray trace

Log Number: 4008219Logbook(s): Entry Makers: beneschBody: 

beta X smaller than design at 8E02?? figure 6 shows something rather different.